OpenStudy (anonymous):

use chain rule to evaluate ((3x)^2-5x)^-2/3

OpenStudy (anonymous):

\[\frac{d}{dx}[(3x)^2-5x)^{-2/3}]= \frac{d}{dx}[(9x^2-5x)^{-2/3}]\] \[ \frac{d}{dx}[(9x^2-5x)^{-2/3}]=\frac{-2(18x-5)}{3(9x^2-5x)^\frac{5}{3}}\]